The length of a rectangle is 5m greater than the width. The perimeter is 150m more than twice the width. Find the width and length of the rectangle. Please help an explain thanks:)!!
Answer by checkley77(12844)

L=W+5
P=2W+150
P=2W+2L
2L=150 REPL;ACE L WITH (W+5) & SOLVE FOR W.
2(W+5)=150
2W+10=150
2W=150-10
2W=140
W=140/2
W=70 ANS. FOR THE WIDTH.
L=70+5=75 ANS FOR THE LENGTH.
P=2*70+2*75
P=140+150
P=290 ANS. FOR THE PERIMETER.